LONG-TERM FINANCING NEEDED

At year-end 2016, total assets for Arrington Inc. were $1.9 million and accounts payable were $450,000. Sales, which in 2016 were $2.8 million, are expected to increase by 30% in 2017. Total assets and accounts payable are proportional to sales, and that relationship will be maintained; that is, they will grow at the same rate as sales. Arrington typically uses no current liabilities other than accounts payable. Common stock amounted to $375,000 in 2016, and retained earnings were $340,000. Arrington plans to sell new common stock in the amount of $180,000. The firm's profit margin on sales is 3%; 55% of earnings will be retained.

  1. What were Arrington's total liabilities in 2016? Write out your answer completely. For example, 25 million should be entered as 25,000,000. Round your answer to the nearest cent.
    $
  2. How much new long-term debt financing will be needed in 2017? Write out your answer completely. For example, 25 million should be entered as 25,000,000. Do not round your intermediate calculations. Round your answer to the nearest cent. (Hint: AFN - New stock = New long-term debt.)
    $
Answer #1

(a) 2016: Total Assets = $ 1.9 million or $ 1900000, Accounts Payable = $ 450000, Common Stock Value = $ 375000, Retained Earnings = $ 340000

Current Liabilities = Accounts Payable = $ 450000 (as the firm uses only accounts payable as current liabilities)

Therefore, Total Liabilities = Total Assets - Common Stock - Retained Earnings = 1900000 - 375000 - 340000 = $ 1185000

(b) Expected Sales Growth in 2017 = 30%, Sales in 2016 = $ 2.8 million

Expected Sales in 2017 = 2800000 x 1.3 = $ 3640000

Total Assets and Accounts Payable are both directly proportional to Sales level.

Hence, Expected Total Assets in 2017 = 1900000 x 1.3 = $ 2470000 and Expected Accounts Payable in 2017 = 450000 x 1.3 = $ 585000

Profit Margin on Sales = 3 % of Sales, Profit of 2017 = 0.03 x 3640000 = $ 109200

Addition to Retained Earnings = 55 % of Profit = 0.55 x 109200 = $ 60060

Long-Term Debt = Total Assets - Common Stock - Retained Earnings - Current Liabilities (Accounts Payable) = 1900000 - 375000 - 340000 - 450000 = $ 735000 ( one can use 2016 values here as long-term debt is not proportional to sales and hence remain the same)

Expected Retained Earnings in 2017 = 2016 Retained Earnings + Addition to Retained Earnings = 340000 + 60060 = $ 400060

Additional Financing Needed (AFN) = Expected Assets in 2017 - Expected Accounts Payable in 2017 - (Common Stock) (this remains constant as the common stock is not proportional to sales volume) - Expected Retained Earnings in 2017 - Constant Long-Term Debt = 2470000 - 585000 - 375000 - 400060 - 735000 = $ 374940

New Stock Issued = $ 180000

Hence, New Long-Term Debt Raised = AFN - New Stock Issued = 374940 - 180000 = $ 194940 (AFN can be raised in two forms, debt and stock. Hence, AFN = Debt to be Raised + Stock to be Issued)